# Assessing Language Comprehension in Large Language Models Using Construction Grammar

### v1

@@ 4.2 Empirical Evaluation and Analysis
We test three OpenAI models333https://platform.openai.com/docs/models on our constructional NLI dataset. We test the GPT-4o-2024-05-13 and GPT-3.5-turbo-0125 models as well as the recently released o1-preview-2024-09-12 model. These models were chosen due to their large sizes making them illustrative examples of the capabilities of state-of-the-art LLMs in general. We also test two LLaMa models, LLaMA-3-8b-instruct and LLaMA-3-70b-instruct444Full results for the LLaMa experiments can be found in Appendix J. We test 3 main scenarios: zero-shot, in-context learning with examples randomly selected from Stanford NLI (SNLI, Bowman et al. 2015), and in-context learning with Constructional NLI (CxNLI, our dataset). We additionally perform Chain-of-Thought (CoT) experiments (Wei et al., 2022b) in each of these scenarios, in an attempt to more precisely explore gaps in model capabilities. Surprisingly, we find that Chain-of-Thought does not lead to performance gains on our dataset.555Full results for CoT are reported in Appendix F A summary of our results for this experiment are reported in Table 2.666We experiment with a variety of prompt formats and report results for the best performing prompts. Details on prompt variations and results can be found in Appendix D
| Setting | IC Data | GPT Accuracy | ||
| 3.5 | 4o | o1* | ||
| Zero-shot | None | 0.6 | 0.88 | 0.90 |
| One-shot | CxNLI | 0.69 | 0.9 | - |
| Three-shot | CxNLI | 0.79 | 0.96 | 0.90 |
| One-shot | SNLI | 0.58 | 0.86 | - |
| Three-shot | SNLI | 0.59 | 0.86 | 0.89 |
Overall, we see that performance is quite high even in the zero-shot setting for GPT-4o and GPT-o1. We also observe that GPT-4o consistently perform better than GPT-3.5. Adding additional examples of Cxns for in-context learning boosts performance, while additional SNLI examples do not boost performance. This is especially true for GPT-3.5, which benefits substantially more from in-context learning from CxNLI. This reliance on in-context learning indicates that our datasets test a different axis of semantic knowledge than more general datasets like SNLI.
5 Experiment 2: Challenge Constructional NLI
@@ 5.1 Dataset
In the previous section, we show that GPT-4o and GPT-3.5 perform impressively on an NLI task that specifically targets constructional semantics. Many of these constructions are common to the English language, and the templates we generate target aspects of meaning that are highly salient for the construction. However, we wish to test LLM generalization of constructional knowledge even in settings with unexpected uses of the construction. In order to do this, we develop a more challenging and NLI dataset (CxNLI-challenge). For 5 of our 8 Cxns, we introduce more challenging examples, which share a surface syntax with our Cxns, but convey a different meaning. To demonstrate this, consider the following two examples:
@@
In the above two examples, it is clear that their syntactic forms appear identical. Both have a subject pronoun, a verb, an object noun phrase followed by an adjective. However, the two Cxns convey different meanings: In Example 5.1 the adjective is the result of the action of the verb, whereas in Example 5.1, the adjective is the state of the noun during the action of the verb, but it is not the resulting state of the action. This difference in meaning is associated with two different Cxns, specifically the resultative and the depictive. We can tease apart this difference in meaning by leveraging our template based hypotheses from our CxNLI dataset in 4. Consider the following templatic hypotheses:
@@
As we can see in Examples 5.1 and 5.1, templatically generating hypotheses for these two examples leads to different relations to the premises. We create this more challenging dataset (CxNLI-Challenge) by gathering "false positives" of Cxns from corpora based on simple pattern matching queries. These "false positive" constructions are those which on the surface appear to be the same as the constructions which we are testing, but have drastically different meanings. We templatically generate hypotheses for these false positives, which produces a dataset very similar in form to CxNLI (described in 4), but with different relations between premises and hypotheses, due to the different meanings of the false positive constructions. In many cases, the resulting hypotheses are rarer phrasings for the challenge Cxns than they are for their intended Cxns from CxNLI. We hypothesize that this feature will make them more difficult for the model, as they are less likely to be highly represented in the pretraining data for the LLM. For humans, interpretation is possible because we are able to generalize the meanings coming from certain slots in the Cxn even in unusual contexts. After the final version of this dataset was created, a second author evaluated the dataset, achieving an IAA of 83% with the original judgements.
@@ 5.2 Empirical Evaluation and Analysis
We utilize the CxNLI-challenge dataset in two settings: as in-context examples for our CxNLI dataset from Section 4, and as an evaluation dataset. We first examine what effect, if any, these challenge examples have on CxNLI performance when they are used as the in-context learning examples for our previous dataset. As we show in Table 3, for GPT-4o, performance using CxNLI-challenge as in-context learning for our CxNLI does not boost or harm performance significantly, and is roughly equivalent to performance in the zero-shot setting. On the other hand, performance on CxNLI drops significantly for GPT-3.5 when it is prompted with CxNLI-challenge. This indicates that GPT-3.5 does not have the same constructional knowledge as GPT-4o, and is easily misled by surface syntactic similarity.777We show for best performing LLM prompts in Table 3. See Appendix D for results for full range of prompts for all experiments. We also experiment with CoT for this dataset, finding that it does not substantially affect performance.
| Setting | IC Data | GPT Accuracy | |
| 3.5 | 4o | ||
| Zero-shot | None | 0.60 | 0.88 |
| One-shot | CxNLI-challenge | 0.38 | 0.89 |
| One-shot | CxNLI-challenge,CxNLI | 0.7 | 0.9 |
@@
We further test if LLMs can perform NLI successfully on these challenge examples themselves. As we can see in Table 4, performance is significantly lower on our challenge NLI dataset in almost every prompt setting. The difference in performance is stark: while these examples are admittedly more unnatural sounding and indeed more difficult for humans, the ceiling of human performance (IAA 83%) is well above current LLM performance, even for GPT-4o and GPT-o1. Again we see a large difference between GPT-4o and GPT-3.5, with GPT-3.5 performing below chance in all task settings. We also see that GPT-o1 performs worse than GPT-4o on this task.
| Setting | IC Data | GPT Accuracy | ||
| 3.5 | 4o | o1 | ||
| Zero-shot | None | 0.26 | 0.58 | 0.46 |
| One-shot | CxNLI | 0.26 | 0.53 | - |
| Three-shot | CxNLI | 0.31 | 0.57 | 0.45 |
| One-shot | SNLI | 0.3 | 0.55 | - |
| Three-shot | SNLI | 0.28 | 0.53 | 0.46 |
6 Experiment 3: Challenge Constructional Reasoning
6.1 Dataset
The above experiment on challenge NLI is only possible for some Cxns. For 5 of our 8 Cxns, we successfully create challenge NLI examples based on shared syntactic forms with other Cxns. However, for 3 of our 8 Cxns (let-alone, comparative-correlative, way-manner), creating a challenge dataset is not straightforwardly possible using form-based “false positives.” A clear example of this is the let-alone Cxn: its syntactic pattern is unique to the Cxn, and there are no examples of this exact syntactic pattern that aren’t instances of the let-alone Cxn. Since we cannot create challenge NLI examples for these 3 Cxns with templates, we must find another method for further investigating the semantic understanding of these Cxns.
In order to tackle this problem, we draw on the semantic experiments from Weissweiler et al. (2022). They lay out a group of 7 tests which are designed to probe the semantic understanding of the comparative-correlative while accounting for a range of potential form-based biases and confounding factors. They find that a set of BERT-style models fail at most of these tests. We extend the tests outlined in Weissweiler et al. (2022) to naturally occurring instances of the comparative-correlative, as opposed to artificial examples used in previous work, and we expand our test set to include similar tests for the let-alone and way-manner Cxns. We also change the task to a Question-Answering (QA) task, as opposed to a Masked Language Modeling (MLM) task to better suit the LLMs we test. In our QA scenario, there are always 3 choices: 2 are based on opposing interpretations of the Cxn, and the third choice is “not enough information,” which the model should select when the Cxn does not provide enough information to answer the question.888Descriptions and examples of every test can be found in Appendix G
While these tests are more specific than the NLI examples above, we believe that this dataset serves a similar purpose to the challenge NLI dataset: to more rigorously probe the meaning of Cxns while controlling for the form as much as possible. After creating this dataset for the let-alone, way-manner and comparative-correlative Cxns, we had an additional author annotate the dataset, achieving IAA of 78%.
6.2 Empirical Evaluation and Analysis
| Cxn | GPT Accuracy | ||
| 3.5 | 4o | o1* | |
| Comparative-Correlative | 0.47 | 0.70 | 0.61 |
| Let-Alone | 0.56 | 0.58 | 0.65 |
| Way-Manner | 0.54 | 0.67 | 0.67 |
[…]
Somewhat surprisingly, we find that LLMs remain unable to solve the tasks of Weissweiler et al. (2022) at high accuracy, even in the Chain-of-Thought setting.999More details on CoT experiments in Appendix F. We find that this low performance is consistent across most of the tests from Weissweiler et al. (2022), although there is wide variation on specific test performance depending on the prompt. Unsurprisingly, we again find that GPT-4o and GPT-o1 outperforms GPT-3.5, though the difference is not as stark as in prior experiments.
@@ 7 Error Analysis
7.1 Constructional NLI and challenge NLI
In §4, we describe our Constructional NLI experiments, which find that GPT-4o and GPT-o1 are extremely proficient at Constructional NLI (CxNLI) while GPT-3.5 lags behind substantially. In §5, we find a similar trend for challenge NLI (CxNLI-challenge), but show that even GPT-4o and GPT-o1 do not perform well on CxNLI-challenge examples. For example, though its performance on the CxNLI resultative is near perfect, it struggles with CxNLI-challenge false positives for the resultative in some cases, as shown below.
@@ 7.1 Constructional NLI and challenge NLI
Here, we investigate if some Cxns are harder for LLMs than others. In Figure 1 we report the accuracy by Cxn in our CxNLI and CxNLI-challenge datasets.101010Accuracies are from the highest performing prompt. In general, the same Cxns are difficult across the CxNLI and CxNLI-challenge datasets. Overall, we see that let-alone and comparative-correlative are the weakest Cxns for GPT-4o, though it is strong across the board with a minimum accuracy of 88%. We show an example of GPT-4o misunderstanding the scale of let-alone in Example 10.
[…]
Hypothesis: Beecher’s reputation as a Man of God was easier to accept than his reputation as a preacher.
Correct Response: Contradiction
Model Response: Entailment
GPT-3.5 is much more variable by Cxn, with a maximum accuracy of 92% for the conative and a minimum of 67% for let-alone.
7.2 Challenge Constructional Reasoning
In §6, we show that GPT-4o, GPT-o1 and GPT-3.5 struggle on a QA formulation of a broad set of challenge Cxn reasoning tasks. After examining the errors across all tests, we highlight that there is significant variation in test-by-test performance based on the input prompt. We highlight two possible prompts here, Prompt 1 and Prompt 4.111111Description of prompt variations and results for all experiments can be found in Appendix D The key difference in these prompts is how they task the model with uncertainty. Since some of the QA tests are designed to be unanswerable given the context, the model is supposed to choose the answer not enough information (always choice C in our QA setup). However, we find that the model is unable to balance between two extremes. In some cases, it answers “not enough information” when interpreting the Cxn is sufficient for providing a likely answer, and in other cases, it makes incorrect assumptions about the relationship between elements of the Cxn and context. We can see this in Figure 2. Prompt 4 includes the line “If there is not enough information, choose C,” and does far better on the tests where there is in fact not enough information, but substantially worse on the tests where the Cxn is enough information to provide an answer. An example of this failure is show in Example 11, where the model misinterprets a nonce adjective as somehow relating to the scale of the Cxn, when in fact there is not enough information to answer the question. As a whole, these results show that in our current experimental setup, we cannot get the models to focus specifically on meanings associated with Cxns as opposed to broader context and world knowledge. This leads to incorrect reasoning in cases where the close interpretation of the Cxn outweighs the other factors needed to determine the correct answer.
[…]
Correct Response: Not enough information
Model Response: Pay More
@@ 8 Discussion
Overall, these experiments show the limits of LLM functional understanding of commonplace English Cxns. Regarding our constructional NLI experiments, our results consistently show that LLMs perform better when they encounter in-context examples similar to the actual task they are required to perform, as opposed to to typical SNLI examples. In fact, an increase in the number of SNLI examples leads to deteriorating performance. This result initially seems counterintuitive, as both types of examples provide the same information about the underlying NLI task. However, this can be explained by viewing LLMs’ use of examples as akin to fine-tuning Dai et al. (2023), and instruction tuning—which aims to help models understand instructions Wei et al. (2022a)—as facilitating them to learn in-context directly on instructions Lu et al. (2024). From this perspective, it becomes clear why examples with the same statistical distribution as the task at hand perform better, despite providing the same task-related information. Regarding Cxn reasoning, we find that prompt variation contributes significantly to overall reasoning over Cxns, though all prompt formulations lead to performance substantially below human levels. Overall, our findings indicate GPT-4o and GPT-o1 have substantially more awareness of Cxns than GPT-3.5 has, which may have implications for the usefulness of these respective models for linguistic research on CxG. However, we find that both GPT-4o and GPT-o1 fail to generalize their constructional knowledge sufficiently to tackle challenge scenarios, on both Constructional Reasoning and NLI challenge tasks.
@@ 9 Conclusions and Future Work
In this work, we have shown where even the latest GPT-4o and GPT-o1 models do not demonstrate functional understanding of Cxns. While GPT-4o and GPT-o1 do perform quite impressively on our constructional NLI task, they fail at both challenge NLI and challenge reasoning scenarios, which require generalization of the constructional semantics to infrequent instantiations of the Cxn. Also, we see that GPT-4o substantially outperforms GPT-3.5 in all settings. GPT-3.5 is especially susceptible to being misled by shallow form-based patterns as opposed to meanings, and in-context learning is especially crucial for GPT-3.5. Overall, these experiments show that the constructional awareness of both GPT-4o and GPT-o1 are far more robust than that of GPT-3.5, but their ability to generalize the meaning of Cxns to new contexts, and thus its “understanding”, still lags substantially behind that of humans.
Thus, our targeted series of experiments demonstrated that LLMs do process constructional semantics up to a point, yet our challenge datasets revealed the breaking point of understanding—where speakers are able to generalize the appropriate semantics to constructional slots filled by pragmatically atypical lexical items, but LLMs are much less proficient at this generalization. Overall, we find that CxG serves as a valuable theoretical lens for probing the functional language understanding of LLMs and comparing this with human linguistic knowledge. Greater contributions to resources such as corpora of Cxns will facilitate empirical data on which constructional understanding can be evaluated with more detail.
@@
Appendix A Construction Grammers

### v2

@@ 4.3 Empirical Evaluation and Analysis
We test three OpenAI models333https://platform.openai.com/docs/models on our constructional NLI dataset: GPT-4o-2024-05-13 and GPT-3.5-turbo-0125, as well as o1-preview-2024-09-12. We also test two Llama models: Llama-3-8B-instruct and Llama-3-70B-instruct. These models were chosen for their large sizes, which make them illustrative examples of the capabilities of state-of-the-art LLMs in general. We test 3 main scenarios: zero-shot, in-context learning with examples randomly selected from Stanford NLI (SNLI, Bowman et al. 2015), and in-context learning with Constructional NLI (CxNLI, our dataset).444We experiment with a variety of prompt formats and report results for the best performing prompts; details can be found in Appendix E. We also perform Chain-of-Thought (CoT) experiments (Wei et al., 2022) in each of these scenarios; it does not lead to performance gains: see Appendix F. A summary of our results for Exp 1 are reported in Table 4.
| Setting | IC | Accuracy | ||||
| Data | GPT | Llama 3 | ||||
| 3.5 | 4o | o1* | 8B | 70B | ||
| 0-shot | None | 0.6 | 0.88 | 0.90 | 0.59 | 0.74 |
| 1-shot | CxNLI | 0.69 | 0.9 | - | 0.65 | 0.84 |
| 3-shot | CxNLI | 0.79 | 0.96 | 0.90 | 0.73 | 0.91 |
| 1-shot | SNLI | 0.58 | 0.86 | - | 0.59 | 0.75 |
| 3-shot | SNLI | 0.59 | 0.86 | 0.89 | 0.58 | 0.74 |
Overall, we see that performance is high even in the zero-shot setting for GPT-4o and GPT-o1. We also observe that GPT-4o and Llama 3 70B consistently perform better than their smaller model counterparts GPT-3.5 and Llama 3 8B. Adding examples of Cxns for in-context learning boosts performance, while additional SNLI examples do not boost performance. This is especially true for GPT-3.5 and Llama 3 8B, which benefit substantially more from in-context learning from CxNLI. This reliance on in-context learning indicates that our datasets test a different axis of semantic knowledge than more general datasets like SNLI.
5 Experiment 2: NLI for Distinguishing Syntactically-Identical Cxns
@@ 5.1 Dataset
In Exp 1, we show that the models perform impressively on an NLI task that specifically targets constructional semantics. The Cxns tested are common to the English language, and the templates we generate target aspects of meaning that are highly salient for the Cxn. In Exp 2, we wish to test whether models can generalize the appropriate constructional semantics for syntactically identical phrasal Cxns that should be ascribed distinct semantics. This enables us to determine if models have a robust capability to attribute and understand constructional semantics, or if this capability might be limited to the more entrenched phrasal Cxns of the language that were tested in Exp 1.
| Exp 1 | Exp 2 | ||
| Cxn | Argument Structure Cxn | Syntactically-Identical New Cxn | Cxn |
| Resultative | I brushed my hair smooth. | A famous emperor buried scholars alive. | Depictive |
| Conative | Marco grabbed at the ladder railing. | Other units exploded at this complex. | Intransitive+"at" |
| Caused-motion | We exported nickel to the United States | I introduced her to my boss. | DitransitiveV+NP+PP |
| Intransitive-motion | The 23 scrambled to the rear of the sub. | They’re listening to the same podcast. | Intransitive+"to" |
| Causative-with | Samsung flooded the market with advertising. | I use a mouse with my left hand. | Transitive+"with" |
Thus, for the 5 argument structure Cxns of our 8 Cxns, we add test instances which share a surface syntax with our Cxns, but convey a different meaning.555We test 5 of 8 Cxns from Exp 1 because 3 Cxns lack syntactically identical counterparts with distinct meanings. Table 5 provides examples of the original Cxns used in Exp 1 and parallel, syntactically identical Cxns tested in Exp 2. Consider the following:
@@
In the above two examples, the syntactic forms are identical: both have a subject pronoun, a verb, an object noun phrase followed by an adjective. However, the two Cxns convey different meanings: In Example 5 the adjective is the result of the action of the verb, whereas in Example 5, the adjective is the state of the noun during the action of the verb, but it is not the resulting state of the action. This difference in meaning is associated with two different Cxns, specifically the resultative and the depictive. We can tease apart this difference in meaning by leveraging our template-based hypotheses from our CxNLI dataset in 4. Consider the following templatic hypotheses:
@@
As we can see in Examples 5 and 5, templatically generating hypotheses for these two examples leads to different relations to the premises. The Cxns we use for this dataset are Intransitive + PPAt, Intransitive + PPto, Ditransitive with NP, PP complements, Transitive + PPwith, and the Depictive.666More examples of each of these Cxns along with NLI tuples are shown in Appendix D, Table 10.
Manual Verification After the final version of this dataset was created, a second author evaluated the dataset, achieving an IAA of 83% with the original judgments. The final Exp 2 “CxNLI-Distinction” dataset totals 99 NLI triples.
5.2 Formalism and Task Design
Let be the target Cxn (e.g., Resultative) and be the syntactically identical distractor Cxn (e.g., Depictive). The Cxns are selected such that the entrenchment of the distractor is lower than that of the target: While both Cxns are realized via the same constructional schema, meaning their syntactic templates (e.g., NP V NP ADJ) are identical (), their underlying semantic interpretations are distinct, .
Our evaluation focuses on premises that are instances of the distractor Cxn . As an illustration, consider one such premise, :
Schema:
Example: “A famous emperor buried China’s scholars alive…”
The core of the experiment is to generate two different hypotheses for this single premise, each designed to probe for the meaning of either the (correct) distractor () or (incorrect) target Cxn ().
1.
An NLI tuple probing the correct (Depictive) meaning:
“China’s scholars were fully alive before being buried."
This creates the NLI tuple: .
2.
An NLI tuple probing the incorrect (Resultative) meaning:
“Burying caused the scholars to become alive.”
This creates the NLI tuple: .
Crucially, this experimental design assesses if models can avoid over-generalizing the meaning of the more entrenched target Cxn, , to a less entrenched distractor Cxn, , that shares the same syntactic schema () but has a distinct semantic interpretation (). By measuring overall accuracy, we test a model’s ability to reject the semantics of the highly entrenched Cxn () when presented with a premise from the less entrenched distractor, while simultaneously accepting the correct meaning (). We use performance on the more straightforward instances in Exp 1 as a baseline; therefore, a significant drop in accuracy on this second task indicates a failure to distinguish Cxns based on their subtle semantic cues, highlighting a key difference from human generalization.
@@ 5.3 Empirical Evaluation and Analysis
We utilize this new Exp 2 dataset as an evaluation dataset to test if LLMs can perform NLI successfully on the new phrasal Cxn examples, where performance requires distinguishing the unique semantics of the syntactically identical Cxn. As we can see in Table 6, performance is significantly lower than our results from Exp 1 in almost every prompt setting and across all models. The difference in performance is stark. While these examples are also slightly more difficult for humans, the ceiling of human performance (IAA 83%) is well above current LLM performance, even for GPT-4o, GPT-o1 and Llama 3 70B. Again we see a large difference between the GPT-4o and the smaller models, and also see that GPT-o1 performs worse than GPT-4o.
We take these results as evidence that the ability of models to abstract over the same syntactic slots and assign the appropriate semantics is limited to the more entrenched Cxns of English, of which our original Exp 1 dataset consisted. By shifting to phrasal Cxns that are syntactically identical but potentially less entrenched, we are essentially testing whether or not models can abstract over less data to arrive at a less statistically likely semantic interpretation. This also highlights that the abstraction process in people clearly goes beyond the syntactic character of the slots alone—people are able to balance their knowledge of lexical semantics with their knowledge of constructional semantics to arrive at the most pragmatically likely interpretation. People’s lexical awareness is also imbued with physical world knowledge; thus, people are abstracting over a distinct set of information than what LLMs have access to.
| Setting | IC | Accuracy | ||||
| Data | GPT | Llama 3 | ||||
| 3.5 | 4o | o1* | 8B | 70B | ||
| 0-shot | None | 0.26 | 0.58 | 0.46 | 0.38 | 0.52 |
| 1-shot | CxNLI | 0.26 | 0.53 | - | 0.29 | 0.60 |
| 3-shot | CxNLI | 0.31 | 0.57 | 0.45 | 0.35 | 0.50 |
| 1-shot | SNLI | 0.3 | 0.55 | - | 0.37 | 0.54 |
| 3-shot | SNLI | 0.28 | 0.53 | 0.46 | 0.39 | 0.57 |
@@ 6 Error Analysis
In Exp 1, we describe our CxNLI experiments, which find that GPT-4o, GPT-o1 and Llama 3 70B are extremely proficient at CxNLI while GPT-3.5 and Llama 3 8B lag behind substantially. In Exp 2, we show that all our tested models do not perform well when tested on our CxNLI-Distinction dataset with five additional, syntactically identical Cxns. For example, though GPT-4o’s performance on the CxNLI resultative is near perfect, it struggles to demonstrate understanding of the syntactically identical depictive:
@@
Here, we investigate if some Cxns are harder for LLMs than others. Among Cxns from Exp 1, let-alone and comparative-correlative are the weakest Cxns for GPT-4o, though it is strong across the board with a minimum accuracy of 88%. GPT-3.5 is much more variable by Cxn, with a maximum accuracy of 92% for the conative and a minimum of 67% for let-alone.
In Figure 1 we report the accuracy by Cxn in our Exp 1 NLI and Exp 2 NLI datasets.777Accuracies are from the highest performing prompt. We only visualize GPT-4o for visual clarity, though trends are similar across models. We see performance is lower for all Cxns we test in Exp 2. This provides evidence supporting our hypothesis, outlined in §3, that the entrenchment of a Cxn contributes to model ability to understand the constructional semantics.
@@ 7 Discussion
Our results show a clear discrepancy in the ability of LLMs to process constructional meaning. This discrepancy highlights a failure in how models generalize from frequent patterns to more nuanced, creative uses of language and has significant implications to the development and use of LLMs
Data, Bias, and the Limits of Scale Our Exp 2 findings challenge the prevailing “more data is better” paradigm Kaplan et al. (2020). This implies a need for new data strategies, such as up-sampling rare-but-important structures or using targeted Ye et al. (2025), adversarial fine-tuning Dong et al. (2021) to correct the biases of the base model.
The Need for Diagnostic Evaluation The significant difference in performance between our two experiments shows how broad-coverage benchmarks can overestimate a model’s true linguistic competence. Our work demonstrates the importance of contrastive, diagnostic benchmarks to test specific, theoretically-grounded phenomena.
A Failure of Causal Reasoning and Its Safety Implications The model’s inability to distinguish a Depictive from a Resultative construction is a failure to reason about causality. This is not a niche linguistic error; it has direct implications for AI safety as misunderstanding the difference between a co-occurring state and a caused outcome could lead to catastrophic errors.
Architectural Limitations and Future Directions The systematic nature of this error across models suggests the issue may be rooted in the architecture of these models. Therefore, the path forward may require exploring novel architectures, such as the addition of long term memory Wang et al. (2023).
@@ 8 Conclusions and Future Work
We have shown where even the latest models do not demonstrate a functional understanding of Cxns: although models can generalize the semantics of entrenched Cxns to creative instantiations, the same models cannot robustly distinguish between syntactically identical Cxns with distinct semantic interpretations. While GPT-4o, GPT-o1 and Llama 3 70B do perform quite impressively on our original constructional NLI task, they fail at the NLI scenario requiring constructional distinction, which requires generalization of the appropriate constructional semantics to syntactically-identical Cxns. Also, we see that GPT-4o substantially outperforms GPT-3.5 in all settings, and in-context learning is especially crucial for GPT-3.5. Overall, these experiments show that the constructional awareness of GPT-4o, GPT-o1 and Llama 3 70B are far more robust than that of GPT-3.5 and Llama 3 8B, but their ability to generalize constructional meaning to both novel instantiations and distinct Cxns still lags substantially behind that of humans.
Thus, our targeted series of experiments demonstrate that LLMs do process constructional semantics up to a point, yet our datasets and experiments reveal the breaking point of understanding—where speakers are able to recognize the appropriate constructional semantics despite both novel instantiations and despite the fact that there are multiple, syntactically identical Cxns that could be candidates for interpreting the phrase at hand. Overall, we find that CxG serves as a valuable theoretical lens for probing the functional language understanding of LLMs with a methodology that tests for linguistic generalization beyond memorization and dependency on pretraining priors. Greater contributions to resources such as corpora of Cxns will facilitate empirical data on which constructional understanding can be evaluated with more detail.
@@
Appendix A Construction Grammars